OpenStudy (3psilon):

f(x) = x^3*g(x) , g(-7) = 2 , g'(-7) = -9 . What is f'(7) ?

OpenStudy (anonymous):

apply multiplication rule for derivatives i dont remember that actual name for it

OpenStudy (anonymous):

use the chain rule to find derivative of f(x)

OpenStudy (anonymous):

"product rule"

OpenStudy (anonymous):

\[f \left( x \right)= x^3 \times g \left( x \right) \] you mean this?

OpenStudy (3psilon):

yes

OpenStudy (anonymous):

\[f \prime \left( x \right)= x^3 \times \left( \frac{ d }{ dx } \right)g \left( x \right) + g \left( x \right) \times \left( \frac{ d }{ dx } \right)x^3\]

OpenStudy (anonymous):

\[f \prime \left( x \right)= x^3 \times g \prime \left( x \right) + g \left( x \right) \times 2x^2\]

OpenStudy (anonymous):

\[f \prime \left( x \right)= x^3 \times g \prime \left( x \right) + g \left( x \right) \times 3x^2\] this one

OpenStudy (anonymous):

now to find \[f \prime \left( 7 \right) =7^3 \times \left( -9 \right) + 2 \times \left( 3 \times7^2 \right)\]

OpenStudy (anonymous):

what is g(7), only know g(-7)